OpenStudy (anonymous):

2f-3 ---- = -11 Please explain how to solve 7

OpenStudy (phi):

first, multiply both sides by 7 can you do that ?

OpenStudy (anonymous):

phil i had this one IN THE BAG :[ go on take it

OpenStudy (anonymous):

@phi -77

OpenStudy (phi):

yes on the right side you get -77 if you multiply the left side by 7 what do you get ? any idea ?

OpenStudy (anonymous):

It doesn't cancel out? 49

OpenStudy (phi):

I mean simplify \[ 7 \cdot \frac{(2f-c)}{7} \]

OpenStudy (phi):

remember you can also write that as \[ \frac{7}{7} \cdot (2f-c) \] what is 7/7 ?

OpenStudy (anonymous):

0?

OpenStudy (phi):

anything divided by itself is 1

OpenStudy (anonymous):

wait, 1?

OpenStudy (anonymous):

o

OpenStudy (phi):

yes 1 so you have \[ 1\cdot (2f-3) = -77\] 1 times anything is the anything. in other words, we can write it as \[ 2f-3= -77\] (some how I change the 3 to a c up above. but it's a 3)

OpenStudy (phi):

2f-3= -77 now write +3 on both sides can you do that ? what do you get ?

OpenStudy (anonymous):

2f=-74

OpenStudy (phi):

last step, divide both sides by 2 notice on the left side you will get 2/2 * f

OpenStudy (anonymous):

f=-37

OpenStudy (anonymous):

right?

OpenStudy (phi):

we can check by "putting in -37" for f in the original problem: \[ \frac{2\cdot -37 -3}{7} = -11 \] and see if if the left side simplifies to -11

OpenStudy (phi):

2*-37 -3 is -74 -3 = -77 and -77/7 is -11 so it looks good.

OpenStudy (anonymous):

yep, thank you!